40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 220 of 341  Not logged in ELOG logo
ID Date Author Type Category Subject
  6107   Mon Dec 12 15:24:21 2011 JenneUpdatePSLPMC and MC were both crappy - now realigned

PMC trans was only ~0.79, where it should be ~0.84 something.  The MC was also not stellar. 

I aligned the beam to the PMC, and am now getting PMC trans 0.837 .

Then I aligned the PSL zigzag to the MC, and got MC Refl down to ~0.6 . 

I then aligned the WFS to the unlocked MC, and the MC Trans QPD to the locked MC.

Things seem good.  MC axis is still in a good place, since we get good michelson fringes at the AS port.

  6106   Mon Dec 12 13:02:08 2011 kiwamuUpdateCDSdaqd restarted

I have restarted the daqd process at 1:01 PM since I have added some new ALS's daq channels.

  6105   Mon Dec 12 11:34:40 2011 Leo SingerSummaryGeneralSome design parameters for a Stewart platform

At the suggestion of Rana and Koji, I have worked out some design parameters for a Stewart platform to be used as a vibration isolation device or as a platform for characterization of suspensions.  I have made some initial guesses about the following design requirements:

  • linear travel: 40 microns peak to peak (based on SOS design requirements in LIGO-T950011)
  • angular travel: 3 mrad peak to peak (based on SOS design requirements in LIGO-T950011)
  • payload mass: 5 kg (wild guess of mass of loaded SOS)
  • payload moment of inertia: 0.01 kg m^2 (wild guess)
  • bandwidth: 500 Hz (suggestion of Rana and Koji: ~kHz)

From these assumptions, I have worked out:

  • peak actuator force: 0.88 kN
  • minimum radius of top platform: 15 cm
  • minimum radius of bottom platform: 30 cm
  • minimum height: 26 cm

The combination of high force, high speed, and ~micron travel limits seems to point to piezoelectric actuators.  PI's model P-225.80 would meet the peak push-pull force requirement, but I have not yet determined if it would meet the bandwidth requirement.  Apparently, typical piezoelectric actuators can exert a greater push force than pull force; wonder if one could use an actuator with a smaller force range than the P-225.80 if the actuator is biased by compression.  (Is this what is meant by a "preloaded" actuator?) 

I have attached a PDF explaining how I worked out the actuator force and platform dimensions.  (I'll try to dice up this PDF and put the contents in the Wiki.)  I also have a plant model in MATLAB with which I have been playing around with control schemes, but I don't think that this is ready to show yet.

Here are some tasks that still remain to be done for this preliminary case study:

  • select sensing technologies: integrated linear encoders and/or strain meters, inertial sensing, optical levers, etc.
  • study joints: Koji and Rana suggest flexures; I need to propose the joint geometry and material
  • study internal modes of the platforms and actuators themselves
  • build noise budget

I'd like to ask for input principally on:

  • appropriateness of my design assumptions
  • piezo actuators currently in use in the lab

 Edit: I also added a Mathematica notebook with the inverse kinematics (mapping from platform state to leg lengths) of the platform. 

 

 

 

 

 

Attachment 1: stewart.pdf
stewart.pdf stewart.pdf stewart.pdf stewart.pdf stewart.pdf stewart.pdf stewart.pdf
Attachment 2: stewart.nb
(* Content-type: application/vnd.wolfram.mathematica *)

(*** Wolfram Notebook File ***)
(* http://www.wolfram.com/nb *)

(* CreatedBy='Mathematica 8.0' *)

(*CacheID: 234*)
(* Internal cache information:
NotebookFileLineBreakTest
... 377 more lines ...
  6104   Mon Dec 12 11:16:02 2011 JenneUpdateRF SystemFoam house on EOM

Foam house installed on EOM a few min ago.  We'll leave it until ~tomorrow, then try out the heater loop.

  6103   Sun Dec 11 17:28:36 2011 kiwamuUpdateGreen Lockingstatus update of the Y arm green lock

Quote from #6102

  + Recent goal : automation of the single arm green lock 

As reported in the previous elog entry #6102, the realtime model and screens have been modified.
Here is a summary about what are new in the realtime model.
 
(What are new ?)
  • The top name of the channels has been changed from GCV to ALS      
    => Although the model name itself is still C1GCV to keep the current relations between other computers.
  • I and Q signals on each sensor.
  • LOCKIN modules to detect the sign of the error signals by shaking suspensions.
  • Offset adjusters, which are combination of a controllable epics value and a low pass filter, to allow a smooth length scan.
  • Input matrix. This branches the input signal to the DOFs as well as the LOCKIN modules.
  • Output matrix to allow some combination of actuation (e.g. DARM, CARM, MCL, etc.,)
  • Output switch to enable/disable any feedbacks to the suspensions
  • Output filters before the suspensions. These filters will be usually flat, but enable us to inject some signals and enable some limiters.
     
    Here is the latest medm screen for the modified realtime controller.
    It gives you the idea of how the latest model works.

 ALSscreens.png

  6102   Sat Dec 10 05:27:43 2011 kiwamuUpdateGreen Lockingstatus update of the Y arm green lock

Status update of the Y arm green lock:

  + Recent goal : automation of the single arm green lock

 

(Things done)

  • Implementation of some realtime LOCKIN modules to detect the sign of the error signals.
  • Modification of the realtime control model to accommodate the I/Q MFD signals, which will be available in the near future. (Of course the model file in the svn has been also updated)
  • Update of the medm screens.
  • Scripting of the auto-lock has been 30 % done.
  • Succeeded in automation of closing the ALS loop. (I have tried several times and no failure was observed so far)

(Things to be done)

  • Scripting a routine to detect the sign of the fine sensor signals.
  • Development of a clever length scan algorhythm.
  • Scripting handing off routines.
  • Implementation of some lock-success binary bits to define the ALS state.
  • Implementation of fail-safes.
  6101   Fri Dec 9 20:03:57 2011 ZachUpdateRF SystemLots of current used in Rich's demod box

D0902745-v5 (probably the AP1053's):

Screen_shot_2011-12-09_at_8.00.54_PM.png

Quote:

Those asymmetric currents are same as what I saw with the table top +/-18V supply. If you really don't like it, there could be an option to disconnect CH3/4 in the box.

In any case, this could be a good long-run test of the demod box, couldn't it?

Quote:

I checked the power regulators on the Rich demod box (according to the schematic, D1000217).  The positive one is LM2941CT, and the negative one is LM2991T.  Both accept input voltage up to +26V or -26V respectively.  So my use of +\- 24V to be regulated down to +\- 15V isn't too crazy.  It's a little crazy, but not too crazy.  They recommend having only a 3V difference between the input and output voltages.  We don't have any 18V or 20V power supplies in the regular LSC power supply rack, so Rana suggested using the 24's. 

When I plug in and turn on the Rich box, the current on the +24V power supply goes up by about 0.8A, and the -24V supply goes up by about 0.3A.  That seems like kind of a lot.  Is that too much?  Do I need to find a better plan that involves +\- 18V?  Thoughts?

For now, the Rich box is off, just in case.

 

 

  6100   Fri Dec 9 17:53:31 2011 DenUpdateAdaptive FilteringC1OAF

[Jenne, Den]

AA filters for witness channels are added to the oaf model. It is working now and the number of memory used is not critical.  NO SYNC is not present any more.

  6099   Fri Dec 9 17:44:45 2011 KojiUpdateRF SystemLots of current used in Rich's demod box

Those asymmetric currents are same as what I saw with the table top +/-18V supply. If you really don't like it, there could be an option to disconnect CH3/4 in the box.

In any case, this could be a good long-run test of the demod box, couldn't it?

Quote:

I checked the power regulators on the Rich demod box (according to the schematic, D1000217).  The positive one is LM2941CT, and the negative one is LM2991T.  Both accept input voltage up to +26V or -26V respectively.  So my use of +\- 24V to be regulated down to +\- 15V isn't too crazy.  It's a little crazy, but not too crazy.  They recommend having only a 3V difference between the input and output voltages.  We don't have any 18V or 20V power supplies in the regular LSC power supply rack, so Rana suggested using the 24's. 

When I plug in and turn on the Rich box, the current on the +24V power supply goes up by about 0.8A, and the -24V supply goes up by about 0.3A.  That seems like kind of a lot.  Is that too much?  Do I need to find a better plan that involves +\- 18V?  Thoughts?

For now, the Rich box is off, just in case.

 

  6098   Fri Dec 9 17:15:29 2011 JenneUpdateIOOMC trans is way down

Quote:

I was looking at the trends from the RAMmon, since I did the EOM alignment yesterday, and wanted to compare them to the MC trans, just to make sure the MC was locked during the times I'm examining.  I was dismayed to discover that the MC has lost its oomph, starting around 11:30 this morning.  Den was the only person in the lab to my knowledge at that time, and he claims that he didn't touch the MC until well after lunch.  As you can see from the 8 hour trend attached, we went from normal ~26000 counts to ~15000, and we're slowly decaying from there.

MC refl looks pretty bad on the camera, particularly in YAW.  Investigations are beginning now....

Edit, ~10min later...  I enabled the WFS (I don't know why they were off...when the MC fell out of lock and relocked itself, the WFS didn't come on), and things went basically back to how they should be.  However, the sans-WFS alignment is still totally crappy, so the PSL beam probably needs to be aligned to the MC.  I don't really want to touch the alignment though without an okay from Kiwamu, so I'll wait for him to come in and confirm that he's happy with the current MC.

 Kiwamu and I discussed, and looking at the AS camera with the PRM and SRM misaligned, but MCWFS engaged, things look good.  This means that it's probably the MC that has drifted, and we want to align the MC back to the PSL beam.

  6097   Fri Dec 9 17:08:41 2011 JenneUpdateRF SystemLots of current used in Rich's demod box

I checked the power regulators on the Rich demod box (according to the schematic, D1000217).  The positive one is LM2941CT, and the negative one is LM2991T.  Both accept input voltage up to +26V or -26V respectively.  So my use of +\- 24V to be regulated down to +\- 15V isn't too crazy.  It's a little crazy, but not too crazy.  They recommend having only a 3V difference between the input and output voltages.  We don't have any 18V or 20V power supplies in the regular LSC power supply rack, so Rana suggested using the 24's. 

When I plug in and turn on the Rich box, the current on the +24V power supply goes up by about 0.8A, and the -24V supply goes up by about 0.3A.  That seems like kind of a lot.  Is that too much?  Do I need to find a better plan that involves +\- 18V?  Thoughts?

For now, the Rich box is off, just in case.

  6096   Fri Dec 9 15:49:24 2011 JenneUpdateIOOMC trans is way down

I was looking at the trends from the RAMmon, since I did the EOM alignment yesterday, and wanted to compare them to the MC trans, just to make sure the MC was locked during the times I'm examining.  I was dismayed to discover that the MC has lost its oomph, starting around 11:30 this morning.  Den was the only person in the lab to my knowledge at that time, and he claims that he didn't touch the MC until well after lunch.  As you can see from the 8 hour trend attached, we went from normal ~26000 counts to ~15000, and we're slowly decaying from there.

MC refl looks pretty bad on the camera, particularly in YAW.  Investigations are beginning now....

Edit, ~10min later...  I enabled the WFS (I don't know why they were off...when the MC fell out of lock and relocked itself, the WFS didn't come on), and things went basically back to how they should be.  However, the sans-WFS alignment is still totally crappy, so the PSL beam probably needs to be aligned to the MC.  I don't really want to touch the alignment though without an okay from Kiwamu, so I'll wait for him to come in and confirm that he's happy with the current MC.

Attachment 1: MCtrans_9Dec2011.png
MCtrans_9Dec2011.png
  6095   Fri Dec 9 15:14:41 2011 DenUpdateCDSrelease 2.4

Alex has created a 2.4 branch of the RCD. Jamie, we can try to compile and install it. As a test a did it for c1oaf, it compiles, installs and runs once variables SITE, IFO, RCD_LIBRARY_PATH are properly defined. As we do not want to run one model at 2.4 code and others at 2.1, I recompiled c1oaf back to 2.1. Jamie, please, let me know when you are ready to upgrade to 2.4 release.

  6094   Fri Dec 9 14:33:16 2011 Alex IvanovUpdateAdaptive FilteringC1OAF

Quote:

I tried to figure out why red NO SYNC label became present in the C1OAF_GDS_TP screen after I added AA filters to the C1OAF model.

C1OAF model contains 8 libraries C1OAF_ADAPT for 8 DOF. I changed C1OAF_ADAPT library to C1OAF_ADAPT_AA library where I added 28 AA filters for 28 witness channels. It turns out that if I use this library for all 8 DOF then I see NO SYNC label, if only for one DOF (MCL) then I see green IOP label. This means that using AA filters for each DOF too much channels of filters are created for online system to operate. I think there is some number inside the code that one can not exceed. Analyzing compilation output after "make c1oaf" I figured out that without using AA filters we have 632 filters and using AA we have 856 filters.

For now I'll use AA filters for MCL only.

 I have a feeling we are not fitting into pre-allocated memory space in the shared memory between the front-end process and the epics process. Filter module data is overwriting some other data and that's why we are not getting a sync light. I suggest we upgrade to 2.4 code first and then we will figure out a way to expand memory areas to fit 856 filters.

  6093   Fri Dec 9 13:28:09 2011 DenUpdateAdaptive FilteringC1OAF

I tried to figure out why red NO SYNC label became present in the C1OAF_GDS_TP screen after I added AA filters to the C1OAF model.

C1OAF model contains 8 libraries C1OAF_ADAPT for 8 DOF. I changed C1OAF_ADAPT library to C1OAF_ADAPT_AA library where I added 28 AA filters for 28 witness channels. It turns out that if I use this library for all 8 DOF then I see NO SYNC label, if only for one DOF (MCL) then I see green IOP label. This means that using AA filters for each DOF too much channels of filters are created for online system to operate. I think there is some number inside the code that one can not exceed. Analyzing compilation output after "make c1oaf" I figured out that without using AA filters we have 632 filters and using AA we have 856 filters.

For now I'll use AA filters for MCL only.

  6092   Thu Dec 8 22:44:55 2011 KojiUpdateRF System4ch demod test result

1) Linearity Test

LO input level was +10dBm. The LO freq was 11MHz and 55MHz for CH1 and CH2 respectively.
The IF frequency was fixed at 10kHz.

The amplitude of the RF input was swept from -50dBm to +15dBm.
Basically I and Q output of CH1 and CH2 was quite linear in this amplitude range.

2) Freqency Response

RF input was fixed at -20dBm and the IF frequency was swept from 1kHz to 1MHz.

The response was flat upto 100kHz, and have sensitivity upto 300kHz.

3) Output noise

Noise floor of the output is ~20nV/rtHz. All of the channels behave in the same way.
1/f start from 100Hz.

Attachment 1: RF_DEMOD_TEST_111208.pdf
RF_DEMOD_TEST_111208.pdf RF_DEMOD_TEST_111208.pdf RF_DEMOD_TEST_111208.pdf
  6091   Thu Dec 8 19:48:23 2011 kiwamuUpdateCDSrestarted c1lsc machine and daqd

Since the c1lsc machine became frozen I restarted the c1lsc machined and daqd.

Then I burtrestored c1lsc, c1ass and c1oaf to this evening. They seem running okay.

  6090   Thu Dec 8 16:54:05 2011 JenneUpdateRF SystemInstallation of new demod box

So far:

* New demod box has been placed in the LSC rack.

* An extra set of +\- 24V has been prepared at the terminal block where all the Crates get their power on the LSC rack.  The grounds were all connected up, but the hot wires weren't.  So there were extra spots available, but none that were pre-wired with my voltages.

     * To do this I turned off all the power supplies in the short rack, since I decided to be a live chicken rather than a dead duck.  After hooking up the new terminal block slots, I turned on the power supplies.

 * Make a power cable to go from the terminal block to the box

Still to do:

 * Connect the spare 55MHz LO and the POP (or POX or POY) unused 11MHz LO to the back of the box

 * Move the PD inputs to the new demod box rather than the borrowed AS11 and POP55 boards

 * Plug the I & Q outs for each freq into some spare ADC channels - must first make sure we have 4 open channels.

 * See if it works.

  6089   Thu Dec 8 14:47:28 2011 JenneUpdateIOOEOM aligned to minimize RAM

Quote:

Monitoring good, but remember that the EOM alignment must be done carefully to minimize the RAM before we can use these trends.

 I temporarily diverted the output of the RAMmon PD to a spectrum analyzer (4195 in Spectrum Analyzer mode), and tweaked the EOM alignment until I minimized the 11 and 55 MHz peaks as much as possible.  It was possible to get each individual peak to disappear into the noise (about -70dBm), but to get both minimized simultaneously I wasn't able to get both down into the noise.  I left the 11MHz at about -55dBm, and the 55MHz at about -60dBm.  If Kiwamu's simulation tells us that one is more significant than the other (55, because we use it for MICH?), then we can decide to favor putting that peak in the noise and sacrifice ~10dB in the other peak. 

When I first plugged the PD into the analyzer, I saw 22MHz and 44MHz (small) peaks, but they went away after the first bit of tweaking.

Before having used the analyzer, I was trying to minimize the demodulated signals via StripTool, but that was a slow process.  The spectrum analyzer was obviously much faster.

The PD has been returned to the regular RAMmon electronics.

Next up: putting in the new demod box that Koji tested last night.

  6088   Thu Dec 8 11:59:53 2011 ZachUpdateRF Systemfringing indeed

Here is a trend of 11 & 55 I&Q, along with the EOM temperature and PSL RMTEMP signals. You can see that there is definitely some fringe-like behavior for monotonic changes in temperature. This is consistent with what I have seen on the gyro table in the past.

Some other notes:

  • The EOM temperature (or at least the sensor temperature) seems to track RMTEMP almost exactly when there is no foam box on the EOM. I have verified that the max-min swing here is the same for both signals (~0.77 K).
  • Something crazy appears to happen at ~10:15, and all the RAM signals get much noisier. Does anyone know what happened at this time (2:15am local)?

We ought to get to the bottom of the fringing. The CTE of LiNbO3 is ~2 ppm/K, so given that the wavelength is on the order of 0.5 K, this is probably not caused by the etalon effect (2ppm/K * 0.5K * ~1cm << 1064nm).

RAM_with_temp_overnight.png

  6087   Thu Dec 8 01:26:31 2011 ZachUpdateRF SystemEOM temp sensor modified

I have modified the EOM temperature sensor circuit for the temperature vs. RAM long-term measurements. The only real change is that the sensor is a 100-kOhm thermistor, instead of a 100-Ohm RTD. These semiconductor thermistors (DigiKey P/N 317-1377-ND) are highly nonlinear and can be much more responsive than RTDs, but this difference is much more noticeable at low temperatures.

Frank had told me that the fractional response of the thermistors was so much higher that I could scale the bridge drive current down by the same factor as the resistance was increasing (i.e., 1 mA -> 1 uA, commensurate with 100 ohms -> 100k) and still see a marked improvement. It turns out that at room temperature for this particular sensor the gain enhancement would only be about ~10x, so I only reduced the drive current to ~10 uA, by INCREASING the drive voltage from ~0.1 V -> 1 V, improving the enhancement to ~100x.

Below is a plot of the real nonlinear response of the thermistor, along with a linear approximation at 298.15 K, which gives a coefficient of ~ -4.67 kOhms/K. The differential bridge output voltage response for the new resistance and current is ~7.5 uV/Ohm 2.5 uV/Ohm, bringing the total temperature response before amplification to ~35 mV/K 11.6 mV/K. Looking at a trend of the FSS_RMTEMP channel over a month, we saw that the maximum PSL table temperature fluctuations were ~2 Kpp, so we aimed the maximize resolution by matching +/- 2 K with +/-10 V at the ADC. This was done by using a gain of ~300 in the AD620 that amplifies the differential bridge output. We found that a gain of ~300 put it pretty close, so the grand total calibration ~ 10.5 V/K 3.5 V/K.

Edit (ZK): I screwed up with calculating the bridge response by a factor of three somehow, so I have stricken and restated the calibrations above

 

 

100k_thermistor_response.png

I took a look at the recently acquired temperature data alongside the RAMmon 11 and 55 signals, and it appears that we're seeing the same sort of fringing effects we usually see, with oscillatory RAM levels for a monotonic change in EOM temperature.The odd bit towards the end is caused by the MC losing lock. 

It is going to be very interesting to find out what causes this fringing.

RAM_with_temp.png

  6086   Thu Dec 8 00:45:13 2011 KojiUpdateRF System4ch demod is ready

I have tested the left 2ch of 4ch demod board.

The left most is for 11MHz, and the next one is for the 55MHz.

  6085   Thu Dec 8 00:18:38 2011 ranaSummaryComputer Scripts / ProgramsBURT restore problems / issues in snapshot scripts

I tried to run the seismic StripTool tonight, which seems liek a simple task. But instead I fell through the rabbit hole again...

The seismic.stp couldn't be run from zita/op340m because zita doesn't have EPICS or MEDM and because the op340m version of StripTool cannot load the new file format in which rossa/pianosa save their files.

Once I got it running by sshing in to rossa, I noticed that the BLRMS trends didn't make any sense. Correctly guessed that this was because all of the BP and LP filters were off. Why? Because of bad BURT, that's why.

As it turns out (if you look through the autoburt logs), several of our FE machines haven't been correctly saving snapshots because of some channel count mismatch between old and new SNAP files. So we are not getting the settings restored correctly for these systems when they get booted. Probably, someone has got to suss out the source of the BURT snap messages; it may be that we have to rehash the snap process after any substantial model rebuild.

For c1pemepics, I did a manual restore from the time when Mirko last ELOG'd that BLRMS was trending OK (Nov 23 @ 3 AM).

Seems like we should also get some kind of auto email warning if the BURTs fail in this way. Otherwise, we'll lose a lot of work if it goes on for weeks.

Bottom line: fix the autoburt so that it doesn't fail after model rebuilds.

  6084   Thu Dec 8 00:04:50 2011 ranaUpdateIOORAM Mon is now being demodulated

Monitoring good, but remember that the EOM alignment must be done carefully to minimize the RAM before we can use these trends.

  6083   Wed Dec 7 20:55:44 2011 Vladimir, DenUpdatedigital noiseMatlab error

Quote:

It would be useful to see some plots so we could figure out exactly what magnitude and phase error correspond to "gross" and "miserable".

To show why Matlab is bad in filtering at small cut-off frequencies we did the same thing in Matab, Octave and R: we've taken the low-pass chebyshev filter of the type 1, order 6, ripple 1 dB, the sampling frequency was 16384 Hz and cut-off frequency varied from 1 to 1000 Hz. Here is the plot for the gain of the zpk model versus to cut-off frequency.

gain_cmp.png

We can see that Matlab's gain shows surprising gains for low cut-off frequencies through for > 100 Hz it is fine. In the next table we compare gain from Foton, Matlab, R and Octave for the same filter. So Foton is also good

freq R_gain matlab_gain octave_gain Foton_gain
1 3.05186270655452e-24 4.8824152e-22 3.05186271e-24 3.05186e-24
10 3.04698335947093e-18 1.8268825e-16 3.04698336e-18 3.04698e-18
100 2.99910841393644e-12 2.9991084e-12 2.99910841e-12 2.99911e-12
1000 2.60247212721439e-06 2.6024721e-06 2.60247213e-06 2.60247e-06
  6082   Wed Dec 7 18:47:36 2011 JenneUpdateRF SystemRAM Mon is now being demodulated

There were 2 open outputs on the splitter in the RAMmon (formerly known as Stochmon) box underneath the BS oplev table.  The input to the splitter comes from the Thorlabs PD that we're using as our RAM monitoring PD.  2 of the outputs go to the RMS detection of 11 and 55 MHz.  Now the other 2 (previously terminated) outputs go over to the LSC rack via SMA cable.  The signal on both channels is ~200mV pk-pk, so -10dBm.  One is plugged into the AS11 demod board (which didn't have a PD input yet), and the other goes to POP55's demod board, so POP55 is not what you think it is for now.

Koji is working on checking out the Rich box, which has 4 demodulators, which we will use eventually.  Right now we're just using the already-plugged-in demod boards so we can start looking at some trends of RAM.  We're going to need to find some channels when we're ready for the switchover.

Zach is nearing completion of the mini-update to the temp sensing system.  Once we have the new more sensitive temp sensor in place, we can have a look-see at the similarities between EOM temperature and RAM levels.

  6081   Wed Dec 7 09:56:14 2011 rana imitation of steveOmnistructureEnvironmentTelephone Map

 This is the fone storree 

we should turn off 3977,3979 and 2351 has no function

 

Edit, JD:  x8925 is at the Visitor desk, and isn't on the map.

  6080   Wed Dec 7 02:55:38 2011 kiwamuUpdateGreen Lockinglocking activity tonight

No real progress.

Probably I spent a bit too much time realigning the beat-note optical path.

 

(what I did)

 - Switched on a power supply which was supposed to give +/- 15V for the broadband beat-note PD.
   The power supply had been somehow turned off.
 - Realigned the beat-note path. When we installed the new EOM mount today, we moved some of the green steering mirrors to make a space.
   So we had to realign the downstream of the beat-note path. After the realignment the DC output of the PD was about 120 mV and the signal level of the beat-note was at -20 dBm.
 - Took noise spectrum of the beat-note with the arm cavity locked by the IR-PDH
    The noise curve was almost the same as before (i.e. unknown high frequency white noise above 20 Hz and some low frequency noise which has structures at 1 and 3 Hz).
-  Closed the ALS loop with the coarse sensor. But I was too lazy to go further more. 

Quote from #6076
Tomorrow I will try :
  (1) Using the fine sensor.
  (2) Noise budgeting with the fine sensor.

  6079   Wed Dec 7 00:48:58 2011 kiwamuUpdateRF SystemRealigned incindent pointing to MC
Actually it was already in a good place.
I just realigned the zig-zag mirrors on the PLS table to bring the entire beam axis a little bit upward.
The WFS servo still seems fine. The input pzt mirrors are still within their range.

Quote from #6077

Next step:  Kiwamu needs to find his happy mode cleaner place, and we'll realign the PSL beam to the MC.  The PSL-MC axes were mismatched pretty badly according to Suresh anyway, so this had to be done no matter what.

  6078   Wed Dec 7 00:11:58 2011 DenUpdateAdaptive FilteringOfflineAF

 I did offline adaptive filtering with yesterday's 3 hours of MC-F and GUR1X data. It turns out that normalized-lms can strongly outperform static Wiener filtering!

offlineaf_psd.png 

This is interesting. It might be something inside MC_F that Wiener static does not see. I think the problem is either with seismometer noise or tilt.

Attachment 2: offlineaf_coh.png
offlineaf_coh.png
  6077   Tue Dec 6 21:37:08 2011 JenneUpdateRF SystemNew EOM mount in place

EOM is remounted on the fancy-pants new mount that I crafted.  EOM is also aligned.  2 green mirrors (the first ones to see the beams coming onto the PSL table from the arm transmissions) had to be moved so I could fit the mount in, since the new mount is bigger than the old one.  I put them back, and approximately realigned them, but didn't do any fine alignment.  This must be done before looking at beatnotes again.

After playing with the EOM, the MC was flashing on higher order modes.  The PSL beam has been realigned to make the MC lock on TEM00, and Suresh helped me center on the WFS and MC2T.

Things look okay for now.  Next step:  Kiwamu needs to find his happy mode cleaner place, and we'll realign the PSL beam to the MC.  The PSL-MC axes were mismatched pretty badly according to Suresh anyway, so this had to be done no matter what.

  6076   Tue Dec 6 02:57:44 2011 kiwamuUpdateGreen Locking1st trial of handing off

I succeeded in handing off the servo from that of the ALS to IR-PDH.

However the handing off was done by the coarse sensor instead of the fine sensor because I somehow kept failing to hand off the sensor from the coarse to the fine one.

The resultant rms in the IR-PDH signal was about a few 100 pm, which was fully dominated by the ADC noise of the coarse sensor.

 

Tomorrow I will try :

  (1) Using the fine sensor.

  (2) Noise budgeting with the fine sensor.

 

Here is the actual time series of the handing off.

YarmALS.png

(Upper left ):  intracavity power.
            As the offset was adjusted the power increased to ~ 0.8. Eventually the power becomes close to the nominal value of 1 after the handing off.
(Lower left) : Frequency of the beat-note.
            After the engagement of the ALS servo, I was scanning the arm length and searching for the resonance by changing the error point of this signal.
(Lower right) : IR-PDH signal.
  6075   Tue Dec 6 00:58:34 2011 DenUpdateWienerFilteringOAF current goal

After reducing the digital noise I did offline Wiener filtering to see how good should be online filter. I looked at the MC_F and GUR1_X and GUR1_Y signals. Here are the results of the filtering. The coherence is plotted for MC_F and GUR1_X signals.

oaf_goal_psd.png     oaf_goal_coh.png

We can see the psd reduction of the MC_F below 1 Hz and at resonances. Below 0.3 Hz some other noises are present in the MC_F. Probably tilt becomes important here.

OAF is ready to be tested. I added AA and AI filters and also a highpass filter at 0.1 Hz. OAF workes, MC stays at lock. I looked at the psd of MC_F and filter output. They are comparable, filter output adapts for MC_F in ~10 minutes but MC_F does not go down too much. Determing the right gain I unlocked the MC, while Kiwamu was measuring something. Sorry about that. I'll continue tomorrow during the daytime.

  6074   Tue Dec 6 00:26:00 2011 kiwamuUpdateLSCALS became robust : UGF = 100 Hz

Eventually the instability in the Y end PDH servo turned out to be some kind of an alignment issue.

After carefully realigning the green beam to the Y arm, the UGF of the ALS loop became able to be at more than 50 Hz.

With this UGF it became able to suppress the arm motion to the ADC noise level (few 100 pm in rms).

Now I am scanning the arm length to look for a TEM00 resonance.

 

(the Story)

I have noticed that the spatial fringe pattern of the reflected green light was very sensitive to the pitch motion of ETMY when the green light was locked to the Y arm.

So I realigned the last two launching mirrors to minimize the reflected light. Indeed the misalignment was mainly in the pitch direction.

I basically translated the beam upward by a couple of mm or so.

The amount of the DC reflection is about 2.4 V when it is unlocked and it is now 0.77 mV when the green light is locked.

Quote from #6072

I guess this could be one of the reasons of the unstable behavior in the Y end PDH lock (#6071). (But still it doesn't fully explain the instability).

  6073   Mon Dec 5 19:38:38 2011 JenneUpdateRF SystemEOM mount getting closer....

I have drilled all the holes necessary, and have tapped all but 4 of the holes for the new EOM mount.  The adapter plate is finished and ready to go (including a 10-min iso sonic bath).  The riser that goes between the table and the 9082 mount is drilled but not yet tapped.

  6072   Mon Dec 5 19:21:55 2011 kiwamuUpdateLSCcoarse beat note signal : ADC limited above 30 Hz

The signal observed by the coarse frequency discriminator was actually dominated by the ADC noise above 30 Hz.

It means that once increasing the UGF more than 30 Hz the servo will feed the ADC noise to the test mass and shake it unnecessarily.

I guess this could be one of the reasons of the unstable behavior in the Y end PDH lock (#6071).

(But still it doesn't fully explain the instability).

 

 To improve the situation I am going to do the following actions:

   (1) Installation of a whitening filter (probably use of SR560s)

   (2) Redesign of the servo filter

 

Here is a brief noise budget of the coarse sensor.

Yarm_ALS_coarse.png

Gray curve: free running noise when no servo is applied

Green curve : in-loop noise when the ALS loop is closed with the coarse frequency-discriminator. The UGF was at 30 Hz.

Red curve : ADC noise of the coarse discriminator

Quote from #6071

 So far I still kept failing to increase the UGF of the ALS servo for some reason (see #6024).

  6071   Mon Dec 5 17:44:41 2011 kiwamuUpdateGeneralmy plan tonight

I am going to try handing off the ALS servo to the IR PDH servo on the Y arm and measure the noise.

 - first I need to investigate why the Y end PDH servo becomes unstable when the ALS is engaged with a high UGF.

 

(some notes)

 So far I still kept failing to increase the UGF of the ALS servo for some reason (see #6024).

Every time when I increased the UGF more then 50 Hz, the Y arm PDH lock became unlocked. It needs an explanation and a solution.

Another thing: During several trials in this evening I found the ETMY_SUSPOS_GAIN had been set to 1, so I reset it to 20, which gives us the damping Q of about 5.

 

(Temperature feedback activated)

 As planed in #6024 I have activated the temperature feedback, so that the PZT control signal is offloaded to the temperature. And it seems working fine.

Currently the gain is set to 0.03, which gives us a time constant of ~30 sec for offloading the control signal.

  6070   Mon Dec 5 10:13:13 2011 JenneUpdateAdaptive FilteringC1OAF

Quote:

I've added filter banks for correction path in the C1OAF model to use AA filters.  I compiled and installed the new version. I runs but does not sync. Probably, I've made a mistake in the some names of epics channels. Leave it for now, figure out tomorrow. If someone needs an old version, it is in the /opt/rtcds/caltech/c1/userapps/trunk/isc/c1/models/c1oaf_BACKUP20111204.mdl file. Corresponding medm screen is in the /opt/rtcds/caltech/c1/userapps/trunk/isc/common/medm/OAF_OVERVIEW.adl file.

 The general rule in the 40m is that if it's not an 'emergency', i.e. something is wrong with the computers and preventing the main locking work, no model recompiling-type activities at nighttime. 

Also, if you do things and recompile, you need to do an svn check-in.  That's where backups are kept.  We don't want to clutter folders with backups anymore.

  6069   Mon Dec 5 09:46:21 2011 ZachMetaphysicsRF SystemRAM diagnosis/suppression plan?

Since no one has made any comments, I will assume that everyone is either 100% satisfied with the outline or they have no interest in the project. Under this assumption, I will make decisions on my own and begin planning the individual steps in more detail.

In particular, I will assume that our goal comprises BOTH characterization of RAM levels and mitigation, and I will try to find the best way that both can be achieved as simultaneously as possible. 

  6068   Mon Dec 5 02:55:30 2011 DenUpdateAdaptive FilteringC1OAF

I've added filter banks for correction path in the C1OAF model to use AA filters.  I compiled and installed the new version. I runs but does not sync. Probably, I've made a mistake in the some names of epics channels. Leave it for now, figure out tomorrow. If someone needs an old version, it is in the /opt/rtcds/caltech/c1/userapps/trunk/isc/c1/models/c1oaf_BACKUP20111204.mdl file. Corresponding medm screen is in the /opt/rtcds/caltech/c1/userapps/trunk/isc/common/medm/OAF_OVERVIEW.adl file.

  6067   Sun Dec 4 23:49:38 2011 DenUpdateIOOWFS

Quote:

Yesterday I locked the MC and left at 8 pm. Analyzing the data I saw that MC was locked all time from 8 pm to 12.30 am when it lost lock. Moreover there was no light on transmition and reflected screens at all. I went to the PSL and saw that no red light comes to the MC from PSL, only green. I took infrared sensos to track the laser light. Then I came back to control room to study the medm diagram of the PSL. Then I came back and saw that the laser beam goes to the MC! I returned to control room and saw light on the MC screens. Does someone do something parallel with me through ssh?

I enabled the auto locker and saw the MC locked for a couple of seconds. After that the WFS were turned on automatically and I saw that the signal of the OSEM local sensors of the MC mirrors began to increase. So the WFS master provides not good feedback signal. I thought that it is due to my recompilation of c1mcs with a fixed if-statement line. And may be if c1mcs workes without digital noise and c1ioo with it then there might occur some mismatches and the signal is corrupted. For this assumption I've recompiled c1mcs back to 1e-20 in the if-statement and so added the digital noise back that I saw in the dtt tools.

However, the problem was still present - WFS feedback signal crashed the MC lock. I open the WFS master window and disabled the output to MC. I can see that the C1:IOO-WFS1_PIT_INMON and other input channels have reasonable values 8 - 20 but the output continues to increase up to 1000000. The output was off so the MC stayed at lock. As for now I turned off WFS so no feedback is applied to MC mirros.

With the help of Suresh we have adjusted optics near PMC and input to the MC on the PSL and in the black box where WFS are. Surprisingly, some optics near WFS was not attached to the table. But these mirrors are not used. One screw was near the hole but not screwed in. This mirror is used. Suresh could also rotate other screws. I thought that they must be attached to the table more rigidly.

Then we found that WFS output matrix is wrong and Suresh recalculated it. After that we've locked the MC using WFS. C1:IOO-MC_RFPD_DCMON is 0.7-0.8. 

We also recompiled and reinstalled C1MCS and C1IOO with fixed if-statement and again saw how MC_F curve moves down. WFS error signals are also improved. But still some more work on output matrix is needed.

  6066   Sun Dec 4 13:56:54 2011 DenUpdateIOOWFS

Yesterday I locked the MC and left at 8 pm. Analyzing the data I saw that MC was locked all time from 8 pm to 12.30 am when it lost lock. Moreover there was no light on transmition and reflected screens at all. I went to the PSL and saw that no red light comes to the MC from PSL, only green. I took infrared sensos to track the laser light. Then I came back to control room to study the medm diagram of the PSL. Then I came back and saw that the laser beam goes to the MC! I returned to control room and saw light on the MC screens. Does someone do something parallel with me through ssh?

I enabled the auto locker and saw the MC locked for a couple of seconds. After that the WFS were turned on automatically and I saw that the signal of the OSEM local sensors of the MC mirrors began to increase. So the WFS master provides not good feedback signal. I thought that it is due to my recompilation of c1mcs with a fixed if-statement line. And may be if c1mcs workes without digital noise and c1ioo with it then there might occur some mismatches and the signal is corrupted. For this assumption I've recompiled c1mcs back to 1e-20 in the if-statement and so added the digital noise back that I saw in the dtt tools.

However, the problem was still present - WFS feedback signal crashed the MC lock. I open the WFS master window and disabled the output to MC. I can see that the C1:IOO-WFS1_PIT_INMON and other input channels have reasonable values 8 - 20 but the output continues to increase up to 1000000. The output was off so the MC stayed at lock. As for now I turned off WFS so no feedback is applied to MC mirros.

  6065   Sat Dec 3 18:29:20 2011 DenUpdateAdaptive Filteringcoherence

I've looked through the coherence between the MC length and seismometers after the if-statement problem was fixed. Coherence improved for all seismometers but is still not 1. It is possible that contribution from X, Y, Z directions split the coherence between them but at ~0.2-03 Hz we do not see much coherence for all these directions.

coh_mcl_seis.pdf

I looked at the coherence between MC2 OSEM signal and MC_F when the AUTO LOCKER is ON and OFF. I thought that we'll ses the same coherence for both regimes as laser is locker to the MC length. However, I figured out the coherence is worse when AUTO LOCKER is ON at frequencies 0.2-0.3 Hz.

sensor_locker.pdf

The first idea that comes to mind is that when feedback to the laser is provided, the pressure to the mirrors from the laser beam is changed.

  6064   Sat Dec 3 16:55:52 2011 DenUpdateIOOdigital noise in MC

I looked once again to the local OSEM sensors and MC length signals. Then I replaced 1e-20 to 1e-50 in the if-statement of the iir_filter function. Here I report about the difference of the signals in question.

First we look at the MC2 OSEM local sensor. In the figure below the psd of the signal is presented in three cases - with a free MC2 mirror without feedback, with a feedback signal and with a feedback signal with corrected if-statement. We can see that without FB the wire resonances are high and dumped when OSEMs are on. However we can see that below 1 Hz the psd of the sensor signal with 1e-20 in the if-statement is higher then psd of the sensor signal from free mirror. FB with 1-50 in the if-statement fixes this problem. 

psd_sensors.jpg

If we take a look on the plot of the coherence between GUR1_X and SENSOR signals we can see that coherence is corrupted when 1e-20 is used in the is-statement and is good when 1e-50 is used.

coh_sensors.jpg

 Next we look at the psd of the MC length. We can see how strongly these curves diverge below 1 Hz. The MC_F signal was also corrupted at higher frequencies.

psd_mcl.jpg

 The coherence between MC_F and GUR1_X is also improved.

coh_mcl.jpg

  6063   Fri Dec 2 20:16:41 2011 DenUpdatedigital noisefirst order transition

In order to verify our theory about coherence corruption in linear systems due to the line

if((new_hist < 1e-20) && (new_hist > -1e-20)) new_hist = new_hist<0 ? -1e-20: 1e-20;  

in the /opt/rtcds/caltech/c1/core/release/src/include/drv/fm10Gen.c in the iir_filter function I've changed -20 to other numbers and watched at the coherence input and output of the digital filter cheby1("LowPass", 3, 0.1, 0.5)cheby1("LowPass", 6, 1, 1.5). The sampling rate was 2K. The frequency responce of the filter presented in this figure.

freqz.pdf

The next plot shows psd and coherence of the signal for different numbers in the if-statement line : 1e-20 , 1e-25, 1e-100.

 trans.pdf

We can see that for present value coherence between input and output signals is small even for low frequencies. The psd of the output signal is also corrupted because at low frequencies it should have the same psd as input signal. For 1e-25 and 1e-100 we can see that coherence is close to 1 at low frequencies so if-statement does not work and we have a first order transition from bad to good filter performance with discontinious jump of coherence.

However, for 1e-25 and 1e-100 data is still corrupted by the round-off error. Lack of coherence for high frequencies can be explained by the fact that dtt tools use single precision for data analysis and output is too small to plot a right coherence. But the coherence is also not precisely 1 for low frequencies. Actually, it is 0.99 for this aggresive filter. We use double precision in the real-time code but still for such kinds of filters round-off error is present. As wrote Daniel Sigg for Cheby filter:  "You need a lot more digits than you may naively suspect. In the 8th order example, the output of each SOS is amplified by ~106. This regardless of the fact that the coefficients are all of order 1. If you require a level of 10-3 attenuation in the stop band, you would have lost 9 digits already. Then, add the fact that you have to do of order 104 subtractions to get from 16kHz to 12Hz, loosing another ~2 digits. On top, the high Q section is probably 10 worse than the others and you lost 12 digits. In a real example this may stack up even worse."

Next we need to figure out what effects does round-off error introduce in the performance of the interferometer.

  6062   Fri Dec 2 17:43:46 2011 ranaUpdatedigital noiseFoton error

It would be useful to see some plots so we could figure out exactly what magnitude and phase error correspond to "gross" and "miserable".

  6061   Thu Dec 1 18:30:39 2011 Vladimir, DenUpdatedigital noiseFoton error

Foton Matlab Error     

We investigated some more the discrepancy between Matlab and Foton numbers. The comparison of cheby1(k, 1, 2*12/16384) was done between versions implemented in Matlab, R and Octave. Filters created by R and Octave agree with Foton.

Also, we found that Matlab has gross precision errors for cutoff frequencies just smaller than used in our fitler, for example cheby1(6, 2*3/16384) fails miserably.

  6060   Thu Dec 1 17:33:18 2011 kiwamuUpdateSUSwatchdogs fixed

The watchdogs' issue has been solved and they are now working fine.

It was just because one of the Sorensens had been off.

The Sorensen is the one supplying +5 V in the 1X5 rack.
This +5 V is actually used as a pull-up-current to properly drive the MAX333As (CMOS analog switch) in the coil drivers (D010001).
So this was it.

Quote from #6010

Tonight we noticed that, in fact, the watchdogs don't work for any of the corner optics (I confirmed that they do work for the ETMs).

  6059   Thu Dec 1 12:27:51 2011 ZachMetaphysicsRF SystemRAM diagnosis/suppression plan?

It seems like there is some confusion---or disagreement---amongst the lab about how to proceed with the RAM work (as Rana mentioned at the TAC meeting, we will henceforth refer to it only as "RAM" and never as "RFAM"; those who refuse to follow this protocol will be taken out back and shot).

I would like to provide a rough outline and then request that people reply with comments, so that we can get a collective picture of how this should work. I have divided this into two sections: 1) Methodology, which is concerned with the overall goal of the testing and the procedure for meeting them, and 2) General Issues, which are broadly important regardless of the chosen methodology.

1. Methodology

There are two broad goals:

  • Characterization of extant RAM
    • Measuring the RAM levels existing in an aLIGO-type interferometer without any suppression systems
    • Modeling to estimate the effect on IFO control and corroboration with measurements where possible
      • DC RAM levels contributing offsets to IFO operating point
      • Quasi-DC RAM levels affecting long term detector tuning (e.g., sensing matrix, MICH -> DARM feedforward, etc.)
      • Audio-frequency RAM contributing noise directly via error point modulation
    • Modeling to scale/adapt results from 40m -> aLIGO
  • Mitigation
    • Developing and assessing systems for suppressing RAM
      • Passive: thermal shielding and isolation
      • Active: EOM temperature control
        • Simple temperature stabilization
        • RAM error signal

The question is: which is our goal? The first, the second, or both? If both, what priority is given to which and can/should they be done in parallel? Also, task distribution.

 

2. General Issues

These are loosely related, so they are in random order:

  • Sensing
    • Temperature
      • What is the priority/urgency of a precision AC-bridge-readout temperature sensor?
      • If priority/urgency is low, what is the priority/urgency of upgrading breadboard controller to protoboard version? The common answer will be "make the protoboard version now", but if the urgency of the final AC sensing is high, it may make sense to focus on finalizing that design (after all, other experiments are waiting on a precision temperature controller, and it is not cost-effective to make many temporary controllers as I have done for the 40m).
      • Sensor noise issues
        • What is the sensor-noise-limited temperature stabilization level?
        • What is our willingness to tolerate the thermal low-passing of the EOM can itself (i.e., what is our sensitivity to gradients)?
        • To answer the above questions, we need to perform stabilization tests with several sensors on the same can, with some in loop (averaged) and some out of loop.
        • If we determine that gradients are a problem, we may need to:
          • Design a casing for outside the EOM (inside the foam box) to make the heating uniform, or
          • We may be able to get away with a more customized heater (instead of heating the can from one side as we do now).
    • Optical RAM
      • Stochmon is a nice diagnostic tool, but do we want something better? In particular, we want to have linear signals about a zero-DC-RAM point, which requires phase
        • Where will this sensor be located?
        • What kind of PD will it be? Broadband? Multi-resonant?
        • What sort of electronics will we need? If we are going to use this as an error signal for controlling the EOM temperature, it is just as important as any other IFO readout, since it may couple into all of them.
          • RF pickup is a BIG ISSUE HERE
          • How will the demodulation phases be selected? It should be possible to take TF measurements in certain misaligned (i.e., non-resonant) conditions and adjust the relative phase between the RAM readouts and standard IFO RF readouts such that they are in phase, but this will require some thinking.
          • Lots more, I'm sure
  • Control
    • Method (overlaps some with methodology portion)
      • What is better, simple temperature stabilization or RAM feeback? (More likely, "how much better is RAM feedback?")
      • If RAM feedback is difficult or impossible to implement effectively (see below), is temperature stabilization good enough?
    • Regime
      • Depending on extant RAM levels and on achievable sensing noise, it will be unwise and/or unnecessary to have a RAM control bandwidth above some relatively low frequency (~sub Hz)
        • Gain where RAM suppression is not needed only injects noise into the system
        • This cutoff frequency is largely determined by the thermal response of the system, but additional filtering will likely be necessary to reduce higher-frequency noise coupling (e.g., nonlinear downconversion of high-frequency signals into heater dissipation, etc.)
    • Efficacy
      • If we do RAM feedback, which signal (i.e. which frequency and quadrature) do we minimize?
        • Do we achieve large common-mode reduction across all RF signals, or is there some differential component?
        • In particular, do we make some or all other control signals noisier by stabilizing/minimizing RAM in one channel?
        • If the answer is yes, can we derive an effective control signal from a linear combination of some or all individual RAM signals?

 

There are probably many other issues I have neglected, so please comment on this rough draft as you see fit!

  6058   Thu Dec 1 11:25:10 2011 steveUpdatePEM40m infrastructure holds up well in strong wind condition

 Santa Anna wind speed was locked around 60 kmph last night on campus. The strongest in 30 years.  The lab hold up well. We did not lose  AC power either.

Threes and windows were blown out and  over on campus.

We have 4 sliding glass windows without "heavy-laser proved" inside protection.

We should plan to upgrade ALL  sliding glass windows with metal protection from the inside.The strongest in 30 years.

Attachment 1: santaannawind.png
santaannawind.png
Attachment 2: wind.png
wind.png
ELOG V3.1.3-